# Die Hard: The On-Off-Cycle of Galaxies on the Star Formation Main Sequence

### v1

@@ 4 Imprints of Star Formation Histories
To test this, we plot in Figure 6 the median radial profile of the stellar age for all galaxies of a given class, normalized by their stellar half-mass radius. Indeed, we find noticeable differences among our galaxy classes. Quenched galaxies (in the two leftmost panels) display very flat radial age profiles with $t_{\mathrm{age}}\approx 10\ \mathrm{Gyr}$, with only a faint trend for late-quenching galaxies towards younger stars in the center. Galaxies that fully evolve along the main sequence have consistently younger stellar populations towards their centers, with the youngest average stellar population in the very center (rightmost panel). Rejuvenating galaxies, especially those that are still on the main sequence at $z\approx 0$, produce a “u”-shaped stellar age profile, suggesting that rejuvenation tends to happen outside the stellar half-mass radii, similar to the results reported for high-redshift rejuvenation by Remus & Kimmig (2025). For lower redshifts, Hao et al. (2024) find in MaNGA data a class of blue spirals with signs of rejuvenation, such as a dip in stellar age going from $0.5$ to $1.5\cdot R_{1/2}$ (their profile type three), which is comparable to what we find here. We note that we find a trend of continuously increasing stellar age with radius in the center only for main-sequence-type galaxies. Caught-type galaxies show a similar age distribution as rejuvenating galaxies. This could indicate a degeneracy of radial age profiles with respect to how sudden and steep the transition from below the main sequence to star-forming is.
@@
Curiously, though the RM-type galaxies are the overall youngest, together with the main-sequence galaxies, they tend toward much lower metallicities, comparable to their older RQ counterparts. It seems thus that low stellar metallicity in massive galaxies is found for two cases: either the galaxy formed and quenched rapidly in the early universe, or its star formation was not continuous, instead involving cycles of gas ejection, quiescence and reaccretion of more pristine, metal-poor gas that thus reduced the overall stellar metallicity. Conversely, remaining constantly on the main sequence requires some degree of gas retention, and thus leads to the overall highest stellar metallicities.
@@
Kimmig et al. (in prep.) show that the chemical evolution of galaxies in the Magneticum Pathfinder simulations consistently begins at high [O/Fe] and low [Fe/H] values beyond the range of Figure 8, with the initial enrichment dominated by type II supernovae and thus $\alpha$ elements. The stellar metallicities of early-quenched galaxies are frozen in after a period of high star formation, leading to a final state with the highest [O/Fe] values since no additional stars with higher iron enrichment are formed. The slower but steady contribution from type Ia supernovae for galaxies that have more time to evolve drives the distribution towards the lower values of [O/Fe] and higher values of [Fe/H] simultaneously. Interestingly, rejuvenating galaxies display the lowest $\alpha$ enrichment, which supports the picture of pristine gas accretion as the driver of rejuvenation.
@@
Interestingly, we find that rejuvenating galaxies (dashed magenta line) lie predominantly at either the low- or high-mass end, which is in part because the overall largest group, the early-quenching galaxies (dark red), concentrates around $10^{10.5}-10^{11}\ \mathrm{M}_{\odot}$. By contrast, late-quenching galaxies (dark blue) tend toward slightly higher stellar masses. However, we note that the “u”-shaped rejuvenator distribution could imply two separate physical mechanisms which dominate. On the right, the increasing rejuvenator fraction with increasing stellar mass may be the result of an increasing number of mergers within their evolutionary history (as the most massive objects have experienced the most mergers), as well as the possibility of cooling from the hot halo (e.g. Fabian & Nulsen 1977; Sarazin 1986; González Villalba et al. 2025).
@@
For the low mass end at $M_{*}\approx 10^{10}M_{\odot}$, we find that quenched-early and quenched-late account for $52\%$, which increases up to $70-80\%$ by $M_{*}\approx 10^{11.25}M_{\odot}$. This is because gas falling into halos with a virial mass greater than $M_{\mathrm{vir}}\gtrsim 10^{12}M_{\odot}$ (galaxy stellar masses of $M_{*}\gtrsim 3\times 10^{10}M_{\odot}$) gets shock-heated to high temperatures where cooling becomes inefficient (e.g. Rees & Ostriker 1977). This leads to the apparent knee in the galaxy luminosity function (Schechter 1976) and stellar mass function (e.g. Bullock & Boylan-Kolchin 2017). Consequently, it is the increase in those groups which leads to the decrease in rejuvenated galaxies.
@@
However, simply becoming quenched does not mean a galaxy may not rejuvenate. Instead, we note that the lower rate of rejuvenated galaxies toward intermediate mass galaxies may be required by construction, as to reach stellar masses of $M_{*}>10^{10.5}M_{\odot}$ requires a high SFR over a long period of time, with simply insufficient time to both quench and rejuvenate. Indeed, the late-quenched galaxies are those which rise the strongest with increasing stellar mass, meaning that they represent galaxies which grew significantly in stellar mass by forming it in-situ, before finally running out of gas (or quenching via a dry merger). Simultaneously, low mass galaxies are more easily quenched as they have more shallow potential wells (e.g. Dekel & Silk 1986; Efstathiou 2000), and may therefore be more susceptible to short-term ejection and reaccretion of gas, and galaxies which end at lower masses must have on average spent a longer time with such shallow potential wells compared to more massive galaxies.
[…]
@@
We note that this is assuming an ex-situ fraction of $0\%$, such that the true mass is likely higher, and also that assuming a Kroupa (2001) IMF would only increase the final mass, due to the higher fraction of low mass stars. Consequently, it is essentially by construction that observed lower mass galaxies at $z=0$ must have undergone some periods of quiescence, or must have formed incredibly late in cosmic time (after the peak of the cosmic SFRD).
@@ 5 Gas Flows fueling Star Formation
We then turn to understand the cause of these different evolutionary histories of galaxies. Stars form from giant molecular clouds (GMCs) surpassing the Jeans criterion (Jeans 1902). While the composition of gas phases plays an important role for star formation, the rates fundamentally depend on the amount of gas available to a galaxy. This reservoir is replenished by gas accretion from the environment and depleted by star formation and outflows. Gas flows and star formation can be directly related through the continuity equation assuming conservation of mass, as defined for example by Bouché et al. (2010) in their equation (3):
@@
Here, $\eta$ quantifies the outflow mass loading by supernovae, and R describes the gas return through stellar winds, each relative to the star formation rate. In case of constant accretion rates and outflows dominated by stellar feedback with constant feedback parameters, the so-called ”Bathtub” model predicts an equilibrium state with a gas mass and star formation rate defined by the accretion rate (e.g. Burkert 2017). Such an equilibrium state is expected to settle in roughly within the depletion time $t_{\mathrm{depl}}\equiv t_{\mathrm{ff}}/\epsilon_{\mathrm{ff}}$. Theoretical models have generally found $\epsilon_{\mathrm{ff}}=0.01$ (Krumholz & McKee 2005; Krumholz & Tan 2007), which is broadly mirrored in observations (Lada et al. 2010; Heiderman et al. 2010; Krumholz et al. 2012; Utomo et al. 2018). For giant molecular clouds with $n_{\mathrm{GMC}}\approx 30\mathrm{cm}^{-3}$, assuming $\mu_{\mathrm{GMC}}=2.34\times 10^{-24}$g (Krumholz et al. 2012), we calculate the free fall time as $t_{\mathrm{ff,GMC}}=\sqrt{3\cdot\pi/(32\cdot\mathrm{G}\cdot\rho_{\mathrm{GMC}})}\approx 8\,$Myr. Correspondingly, $t_{\mathrm{dep,GMC}}\approx 0.8\,$Gyr, which may be shorter for denser clouds.
On the other end, for full galaxies at $z=0$ the depletion timescale is longer at $t_{\mathrm{dep,gal}}\approx 2\,$Gyr (Utomo et al. 2018), with the IGM generally more diffuse than the denser GMCs. However, this changes across redshift, as high redshift galaxies are much more gas rich and therefore closer to the limit of GMCs. Thus, we choose as our smoothing window for star formation rates a value in-between of $t_{\mathrm{smooth}}=1\,$Gyr.
We measure in- and outflow rates from particle data in the simulation and derive expected star formation rates according to Equation 1. Flow rates are derived within spherical shells with a thickness of $\Delta r=0.01\times r_{\mathrm{vir}}$. Each SPH particle contributes to flows weighted by its mass, velocity and fraction of volume overlapping with the shell. Similar to the work by Mitchell et al. (2020), we require particles to have a radial velocity of $|v_{\mathrm{rad}}|<50\ \mathrm{km/s}$ to contribute to gas flow rates. For this work, we do not differentiate between outflow by stellar or AGN feedback. Instead, we set $\eta=0$ and use the net accretion rate $\dot{M}_{\mathrm{net}}\equiv\dot{M}_{\mathrm{in}}-\dot{M}_{\mathrm{out}}$ for gas accretion. The return fraction is assumed constant as $\mathrm{R}=0.4$ (Bruzual & Charlot 2003). With this implementation we can derive the evolution of the gas reservoir and respective star formation rates purely from gas accretion rates. Further, we solve the differential equation numerically in order to account for changes in accretion rates and derive gas masses as
@@
We compare the resulting measured inflow rates to expectations from dark matter assembly and smooth accretion according to Dekel et al. (2009) by normalizing our measurements by their halo-mass-dependent and redshift-dependent results. We find that our measured absolute inflow rates are higher by a factor of roughly two to five at the virial radii, with net accretion rates (inflow minus outflow) lying at similar values to the model expectation, both at the virial radius as well as onto the galaxy itself – see Figure 13 in the Appendix, as well as Seidel et al. (to be submitted). Deviations toward higher or lower net accretion are to be expected, as there must be additional gas mixing and ejection that go beyond simpler dark matter inflow assumptions.
@@
Early-quenching galaxies exceed at funneling gas into the center in the early universe, while at later times not even all of inflowing gas onto the halo makes it into the galaxy. The high peak at early times (both in absolute flow rate, see Figure 13, as well as significant mixing in the halo itself shown in Figure 10) may indicate a star burst at early times, which would both rapidly deplete as well as eject the gas, similar to what is shown by Kimmig et al. (2025).
@@
In order to investigate how well the evolution of star formation rates is determined by gas accretion, we show in Figure 11 the agreement between the expected star formation obtained from the “Bathtub” model using the measured gas accretion rates, and the real simulation output for the total galaxy sample across cosmic times. Note that we set the mass loading factor $\eta$ to zero, as we directly measure the outflow rates, accounting therefore for both the contributions from AGN feedback and stellar feedback. The violins indicate the fractions of galaxies that have star formation rates in agreement with expected values when using Equation 1. We find that the fraction of bathtub-like galaxies in the total galaxy sample (solid blue line) quickly rise and peak soon after $z=5$. This means that at high redshifts, the assumed constant values for the return fraction $R$ and depletion time $t_{\mathrm{depl}}$ of a dense giant molecular cloud in a galaxy with a star formation efficiency of $\epsilon_{\mathrm{ff}}=0.01$ describe the transformation of accreted gas into stars well. Going toward lower redshift, then, the agreement drops. We note that there are more elaborate descriptions of the depletion time scale, like a redshift-dependent description following the trend and offset to the main sequence $\Delta\mathrm{MS}$ shown by Tacconi et al. (2020), yet it is noteworthy that our simple approach works well at high redshift.
@@
Figure 11: Violin plot indicating the mean agreement of expectations from net accretion and the star formation rates derived from Equation 1 with simulation output. An agreement is registered when both values deviate less than 0.5 dex from each other or for quenched galaxies, when simulation output and derived values return any SFR below $10^{-3}\mathrm{M}_{\odot}/\mathrm{yr}$. The vertical histograms (violins) count the agreement fraction of each galaxy within $500\ \mathrm{Myr}$ before and after each snapshot, combining three to four snapshot data points. This produces discretized fractions accumulating at 0.0, 0.25, 0.3, 0.5, 0.7, 0.75 and 1. The solid line follows the mean values within each bin. The blue dashed line shows the evolution of the star formation rate density of the galaxy sample normalized to the range of the figure. The black dashed line shows observational results for the evolution of the star formation rate density compiled by Madau & Dickinson (2014) and was normalized as well.
@@
The shape of this agreement is reminiscent of another well known relation. Thus, for comparison, we plot in arbitrary units (as we care about the shape and peak) the star formation rate density (SFRD) of the sample (dashed blue line), which peaks just before $z=2$, in good agreement from the moment of SFRD peak according to observations (Madau & Dickinson 2014). We find a similar form, and slightly earlier peak, which may indicate that one of the reasons for the observed cosmic SFRD is the efficiency at which present gas can be converted into stars, with high redshift galaxies following the bathtub model of direct conversion on short depletion times, while lower redshift galaxies grow increasingly inefficient.

### v2

@@ 4 Imprints of star formation histories
To test this, we plot in Figure 6 the median radial profile of the stellar age for all galaxies of a given class, normalized by their stellar half-mass radius. Indeed, we find noticeable differences among our galaxy classes. Quenched galaxies (in the two leftmost panels) display very flat radial age profiles with $t_{\mathrm{age}}\approx 10\ \mathrm{Gyr}$, with only a faint trend for late-quenching galaxies towards younger stars in the center. Galaxies that fully evolve along the main sequence have consistently younger stellar populations towards their centers, with the youngest average stellar population in the very center (rightmost panel). Rejuvenating galaxies, especially those that are still on the main sequence at $z\approx 0$, produce a “u”-shaped stellar age profile, suggesting that rejuvenation tends to happen outside the stellar half-mass radii, similar to the results reported for high-redshift rejuvenation by remus:2023. For lower redshifts, hao:2024 find in MaNGA data a class of blue spirals with signs of rejuvenation, such as a dip in stellar age going from $0.5$ to $1.5\cdot R_{1/2}$ (their profile type three), which is comparable to what we find here. We note that we find a trend of continuously increasing stellar age with radius in the center only for main-sequence-type galaxies. Caught-type galaxies show a similar age distribution as rejuvenating galaxies. This could indicate a degeneracy of radial age profiles with respect to how sudden and steep the transition from below the main sequence to star-forming is.
@@
Curiously, though the RM-type galaxies are the overall youngest, together with the main-sequence galaxies, they tend toward much lower metallicities, comparable to their older RQ counterparts. It seems thus that low stellar metallicity in massive galaxies is found for two cases: either the galaxy formed and quenched rapidly in the early universe, or its star formation was not continuous, instead involving cycles of gas ejection, quiescence and re-accretion of more pristine, metal-poor gas that thus reduced the overall stellar metallicity. Conversely, remaining constantly on the main sequence requires some degree of gas retention, and thus leads to the overall highest stellar metallicities.
@@
Kimmig et al. (in prep.) show that the chemical evolution of galaxies in the Magneticum Pathfinder simulations consistently begins at high [O/Fe] and low [Fe/H] values beyond the range of Figure 8, with the initial enrichment dominated by type II supernovae and thus $\alpha$ elements. The stellar metallicities of early-quenched galaxies are frozen in after a period of high star formation, leading to a final state with the highest [O/Fe] values since no additional stars with higher iron enrichment are formed. The slower but steady contribution from type Ia supernovae for galaxies that have more time to evolve drives the distribution towards the lower values of [O/Fe] and higher values of [Fe/H] simultaneously. Interestingly, rejuvenating galaxies display the lowest $\alpha$ enrichment, which supports the picture of pristine gas accretion as the driver of rejuvenation. The detailed treatment considering mass-dependent lifetime functions and metallicity-dependent stellar yields of AGB stars, SNII and SNIa, produce metallicities consistent with observations (dolag:2017; kudritzki:2021; dolag:2025). Since the dust and gas phases do not segregate on the galactic scale, we expect uncertainties in the IMF and stellar yields models (including for example the binary fraction) to dominate over the impact of dust. Nonetheless, the values shown in this work are to be compared to dust-corrected results.
@@
Interestingly, we find that rejuvenating galaxies (dashed magenta line) lie predominantly at either the low- or high-mass end, which is in part because the overall largest group, the early-quenching galaxies (dark red), concentrates around $10^{10.5}-10^{11}\ \mathrm{M}_{\odot}$. By contrast, late-quenching galaxies (dark blue) tend toward slightly higher stellar masses. However, we note that the “u”-shaped rejuvenator distribution could imply two separate physical mechanisms which dominate. On the right, the increasing rejuvenator fraction with increasing stellar mass may be the result of an increasing number of mergers within their evolutionary history (as the most massive objects have experienced the most mergers), as well as the possibility of cooling from the hot halo (e.g. fabian:1977:nulsen; sarazin:1986; gonzalez_villalba:2025). To investigate the cause of the bimodal distribution, splitting the rejuvenators up by their stellar mass into high- and low-mass subclasses should reveal collective differences in gas accretion evolution, in-situ fractions and merger histories. The latter would be a plausible main driver for rejuvenation of the higher-mass galaxies for which smooth accretion is quenched by shock heating by their hot halos (e.g. dekel:2006:birnboim). Furthermore, explicitly tracing gas particles would provide further insight into the different mass assembly histories, which is a topic for future work.
@@
For the low mass end at $M_{*}\approx 10^{10}M_{\odot}$, we find that quenched-early and quenched-late account for $52\%$, which increases up to $70-80\%$ by $M_{*}\approx 10^{11.25}M_{\odot}$. This is because gas falling into halos with a virial mass greater than $M_{\mathrm{vir}}\gtrsim 10^{12}M_{\odot}$ (galaxy stellar masses of $M_{*}\gtrsim 3\times 10^{10}M_{\odot}$) gets shock-heated to high temperatures where cooling becomes inefficient (e.g. rees:1977:ostriker). This leads to the apparent knee in the galaxy luminosity function (schechter:1976) and stellar mass function (e.g. bullock:2017:boylan-kolchin). Consequently, it is the increase in those groups which leads to the decrease in rejuvenated galaxies.
@@
However, simply becoming quenched does not mean a galaxy may not rejuvenate. Instead, we note that the lower rate of rejuvenated galaxies toward intermediate mass galaxies may be required by construction, as to reach stellar masses of $M_{*}>10^{10.5}M_{\odot}$ requires a high SFR over a long period of time, with simply insufficient time to both quench and rejuvenate. Indeed, the late-quenched galaxies are those which rise the strongest with increasing stellar mass, meaning that they represent galaxies which grew significantly in stellar mass by forming it in-situ, before finally running out of gas (or quenching via a dry merger). Simultaneously, low mass galaxies are more easily quenched as they have more shallow potential wells (e.g. dekel:1986:silk; efstathiou:2000), and may therefore be more susceptible to short-term ejection and re-accretion of gas, and galaxies which end at lower masses must have on average spent a longer time with such shallow potential wells compared to more massive galaxies.
Generally, we note that rejuvenated galaxies exist at all stellar masses, while our main-sequence galaxies (cyan), though few in number, show a clear trend toward higher $z=0$ masses. This is a direct consequence from the constraint of lying on or near the main sequence at all times. When taking an initial galaxy of $M_{*}=10^{8}M_{\odot}$ at $z=4,3,$ and $2$, assuming a Chabrier IMF (chabrier:2003) with mass loss following leitner:2011 and determining its star formation rate in time steps of 1010\,Myr based on the evolving pearson:2018 main sequence, we determine the expected final $z=0$ mass as $M_{*}=6\times 10^{10}M_{\odot},\,4\times 10^{10}M_{\odot}$ and $3\times 10^{10}M_{\odot}$.
@@
We note that this is assuming an ex-situ fraction of $0\%$, such that the true mass is likely higher, and also that assuming a kroupa:2001 IMF would only increase the final mass, due to the higher fraction of low mass stars. Consequently, it is essentially by construction that observed lower mass galaxies at $z=0$ must have undergone some periods of quiescence, or must have formed incredibly late in cosmic time (after the peak of the cosmic SFRD).
@@ 5 Gas flows fueling star formation
We then turn to understand the cause of these different evolutionary histories of galaxies. Stars form from giant molecular clouds (GMCs) surpassing the Jeans criterion (jeans:1902). While the composition of gas phases plays an important role for star formation, the rates fundamentally depend on the amount of gas available to a galaxy. This reservoir is built up and replenished by gas accretion from the environment and depleted by star formation and outflows. Gas flows and star formation can be directly related through the continuity equation assuming conservation of mass, as defined for example by bouche:2010 in their equation (3):
@@
Here, $\eta$ quantifies the outflow mass loading by supernovae, and R describes the gas return through stellar winds, each relative to the star formation rate. In case of constant accretion rates and outflows dominated by stellar feedback with constant feedback parameters the “Bathtub” model predicts an equilibrium state with a gas mass and star formation rate proportional to the accretion rate (e.g. burkert:2017) when star formation and feedback cancel out gas accretion. This final galaxy gas reservoir determined by a constant accretion rate serves as the volume of the figurative bathtub. Such an equilibrium state is expected to settle in roughly within the depletion time $t_{\mathrm{depl}}\equiv t_{\mathrm{ff}}/\epsilon_{\mathrm{ff}}$. Theoretical models have generally found $\epsilon_{\mathrm{ff}}=0.01$ (krumholz:2005; krumholz:2007), which is broadly mirrored in observations (lada:2010; heiderman:2010; krumholz:2012; utomo:2018). For giant molecular clouds with $n_{\mathrm{GMC}}\approx 30\mathrm{cm}^{-3}$, assuming a mean molecular mass $\mu_{\mathrm{GMC}}=2.34\times 10^{-24}$g (krumholz:2012), we calculate the free fall time as $t_{\mathrm{ff,GMC}}=\sqrt{3\cdot\pi/(32\cdot\mathrm{G}\cdot\rho_{\mathrm{GMC}})}\approx 8\,$Myr. Correspondingly, $t_{\mathrm{dep,GMC}}\approx 0.8\,$Gyr, which may be shorter for denser clouds.
On the other end, for full galaxies at $z=0$ the depletion timescale is longer at $t_{\mathrm{dep,gal}}\approx 2\,$Gyr (utomo:2018), with the IGM generally more diffuse than the denser GMCs. However, this changes across redshift, as high redshift galaxies are much more gas rich and therefore closer to the limit of GMCs. Thus, we choose as our smoothing window for star formation rates a value in-between of $t_{\mathrm{smooth}}=1\,$Gyr.
We measure in- and outflow rates of galaxies from particle data in the simulation and derive expected star formation rates according to Equation 1. The flow rates are derived by drawing a thin spherical shell around each galaxy at a radius of $r_{\mathrm{gal}}=0.1\times r_{\mathrm{vir}}$ and a thickness of $\Delta r=0.01\times r_{\mathrm{vir}}$. Each SPH gas particle contributes to flow rates weighted by its mass, velocity and fraction of its volume overlapping with the shell volume around the respective galaxy. Similar to the work by mitchell:2020, we require particles to have a radial velocity of $|v_{\mathrm{rad}}|<50\ \mathrm{km/s}$ to contribute to gas flow rates. For this work, we do not differentiate between outflow by stellar or AGN feedback. Instead, we set $\eta=0$ and use the net accretion rate $\dot{M}_{\mathrm{net}}\equiv\dot{M}_{\mathrm{in}}-\dot{M}_{\mathrm{out}}$ for gas accretion. The return fraction is assumed constant as $\mathrm{R}=0.4$ (bruzual:2003:charlot). With this implementation we can derive the evolution of the gas reservoir and respective star formation rates purely from gas accretion rates. Further, we solve the differential equation numerically in order to account for changes in accretion rates and derive gas masses as
@@
We compare the resulting measured inflow rates to expectations from dark matter assembly and smooth accretion according to dekel:2009 by normalizing our measurements by their halo-mass-dependent and redshift-dependent results. We find that our measured absolute inflow rates are higher by a factor of roughly two to five at the virial radii, with net accretion rates (inflow minus outflow) lying at similar values to the model expectation, both at the virial radius as well as onto the galaxy itself – see Figure 13 in the Appendix, as well as Seidel et al. (in prep.). Deviations toward higher or lower net accretion are to be expected, as there must be additional gas mixing and ejection that go beyond simpler dark matter inflow assumptions.
@@
Early-quenching galaxies exceed at funneling gas into the center in the early universe, while at later times not even all of inflowing gas onto the halo makes it into the galaxy. The high peak at early times (both in absolute flow rate, see Figure 13, as well as significant mixing in the halo itself shown in Figure 10) may indicate a star burst at early times, which would both rapidly deplete as well as eject the gas, similar to what is shown by kimmig:2025a.
@@
In order to investigate how well the evolution of star formation rates is determined by gas accretion, we show in Figure 11 the agreement between the expected star formation obtained from the “Bathtub” model using the measured gas accretion rates, and the real simulation output for the total galaxy sample across cosmic times. Note that we set the mass loading factor $\eta$ to zero, as we directly measure the outflow rates, accounting therefore for both the contributions from AGN feedback and stellar feedback. The violins indicate the fractions of galaxies that have star formation rates in agreement with expected values when using Equation 1. We find that the fraction of “Bathtub”-like galaxies in the total galaxy sample (solid blue line) quickly rise and peak soon after $z=5$. This means that at high redshifts, the assumed constant values for the return fraction $R$ and depletion time $t_{\mathrm{depl}}$ of a dense giant molecular cloud in a galaxy with a star formation efficiency of $\epsilon_{\mathrm{ff}}=0.01$ describe the transformation of accreted gas into stars well. Going toward lower redshift, then, the agreement drops. We note that there are more elaborate descriptions of the depletion time scale, like a redshift-dependent description following the trend and offset to the main sequence $\Delta\mathrm{MS}$ shown by tacconi:2020, yet it is noteworthy that our simple approach works well at high redshift.
@@
Figure 11: Violin plot indicating the mean agreement of expectations from net accretion and the star formation rates derived from Equation 1 with simulation output. An agreement is registered when both values deviate less than 0.5 dex from each other or for quenched galaxies, when simulation output and derived values return any SFR below $10^{-3}\mathrm{M}_{\odot}/\mathrm{yr}$. The vertical histograms (violins) count the agreement fraction of each galaxy within 500Myr500\ $\mathrm${Myr} before and after each snapshot, combining three to four snapshot data points. This produces discretized fractions accumulating at 0.0, 0.25, 0.3, 0.5, 0.7, 0.75 and 1. The solid line follows the mean values within each bin. The blue dashed line shows the evolution of the star formation rate density of the galaxy sample normalized to the range of the figure. The black dashed line shows observational results for the evolution of the star formation rate density compiled by madau:2014:dickinson and was normalized as well.
@@
The shape of this agreement is reminiscent of another well known relation. Thus, for comparison, we plot in arbitrary units (as we care about the shape and peak) the star formation rate density (SFRD) of the sample (dashed blue line), which peaks just before $z=2$, in good agreement from the moment of SFRD peak according to observations (madau:2014:dickinson). We find a similar form, and slightly earlier peak, which may indicate that one of the reasons for the observed cosmic SFRD is the efficiency at which present gas can be converted into stars, with high redshift galaxies following the “Bathtub” model of direct conversion on short depletion times, while lower redshift galaxies grow increasingly inefficient. This is consistent with previous findings by teklu:2023 where the overall growth in size leads to lower star formation rates with the same amount of cold gas available. In future work, we plan to more directly explore the link between the internal gas distribution of galaxies, a detailed characterization of their accretion flows (e.g. smoothness and angular momentum), and the resultant agreement with the “Bathtub” model.